OpenStudy (anonymous):

ind all real numbers that satisfy each equation: cos(x/2) = 1/2

OpenStudy (jdoe0001):

$$ \bf cos(\frac{x}{2}) = \frac{1}{2}\\ cos^{-1}\left(cos\left(\frac{x}{2}\right)\right) = cos^{-1}\left(\frac{1}{2}\right) $$

OpenStudy (anonymous):

\[\frac{ x }{ 2 }=\frac{ \pi }{ 3 }+2n \pi\] or \[\frac{ x }{2 }=\frac{- \pi }{3 }+2n \pi\] so:\[x=\frac{ 12n \pi+2\pi }{ 3 }\] or \[x=\frac{ 12n \pi-2\pi }{ 3 }\]

OpenStudy (anonymous):

how do you know if you subtract or add the 2pi or pi

OpenStudy (anonymous):

and the answer says 4pi is that just cause its not restricted?

OpenStudy (anonymous):

sin & cos with period 2π

OpenStudy (anonymous):

n=0,1,2,3,.................... if n=0 ,x=pi/3 or x=-pi/3 if n=1 ,x=7pi/3 or x=-7pi/3 and so on.